Area of a painting. A rectangular painting with a width of x centimeters has an area of x^2 + 50x square centimeters. Find a binomial that represents the length.
Answer by ptaylor(2198)


Area of rectangle(A)=Length(L)*Width(W) or A=L*W
We are told that:
A=x^2+50x sq cm, and
W=x cm
So, our equation to solve is:
x^2+50x=L*x divide each term by x
x+50=L or
L=x+50 cm------------------------------Length
CK
A=L*W=(x+50)*x=x^2+50x
Hope this helps---ptaylor